OpenStudy (trisarahtops):

Where is the second derivative of y = 2xe−x equal to 0? 0 1 2 4

OpenStudy (trisarahtops):

@ParthKohli

zepdrix (zepdrix):

\[\large\rm y=2xe^{-x}\]Do you understand how to find the first derivative? :)

OpenStudy (solomonzelman):

Use the product rule: d/dx (f•g) = f'•g + f•g'

OpenStudy (solomonzelman):

Do you want an example (with some notes and explanations)?

OpenStudy (trisarahtops):

I got -2e^-x(x-1)

zepdrix (zepdrix):

For first derivative? Looks good so far :)

OpenStudy (solomonzelman):

the first, that is right...

OpenStudy (trisarahtops):

okay good. now what?

OpenStudy (solomonzelman):

second derivative

OpenStudy (solomonzelman):

differentiate the result (that you obtained for the first derivative). This wll give you the 2nd derivative.

OpenStudy (trisarahtops):

(-2e^-x(x-3) ??

OpenStudy (trisarahtops):

oh no didn't you say it would be 2e^-x(x-2)

OpenStudy (solomonzelman):

yes, checked it, correct

OpenStudy (solomonzelman):

but you should differentiate it yourself, instead of saying; wouldn't you want to do it yourself (tho)?

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle y^{(0)}(x)=kxe^{bx} }\) \(\large\color{#000000 }{ \displaystyle y^{(1)}(x)=ke^{bx}+bkxe^{bx} }\) \(\large\color{#000000 }{ \displaystyle y^{(2)}(x)=ke^{bx}+bke^{bx}+b^2kxe^{bx} }\) \(\large\color{#000000 }{ \displaystyle y^{(2)}(x)=(k+bk)e^{bx}+b^2kxe^{bx} }\)

OpenStudy (solomonzelman):

in general, (with chain rule to the exponent, and the product rule of course)

OpenStudy (solomonzelman):

In any case, you have to set f''(x)=0 and solve for x.

OpenStudy (solomonzelman):

2e^-x(x-2)=0

OpenStudy (solomonzelman):

2e^(-x)\(\ne\)0 \(\forall\)x

OpenStudy (solomonzelman):

So, (x-2)=0 good luck!